OpenStudy (anonymous):

the surface of soccer ball is made of a pattern of regular pentagons and hexagons .If each hexagon on a soccer ball has perimeter of 9 inches , what is the area of a hexagon?

OpenStudy (anonymous):

Hexagon is made up of 6 equilateral triangles. Since they are equilateral and the perimeter is 9 inches, then each side must be 3 Area of a triangle is 1/2 b h we need to find the height by drawing the height of an equilateral it would create a right triangle the hypotenuse would be 3 the bottom leg would be 1 1/2 to find the height you use the pythagorean theorem a^2 + b^2 = c^2 (3/2)^2 + b^2 = 3^2 9/4 + b^2 = 9 b^2 = 6 3/4 = 27/4 b = 3 sqrt(3) over 2 Area would be 1/2 * 3 * 3sqrt(3)/2 9sqrt(3) ------- 4 then you have 6 of these triangles. 6 * 9 sqrt(3) -------- 4 27 sqrt(3) -------- 3

OpenStudy (anonymous):

thank